# Dissipating quartets of excitations in a superconducting circuit

### v1

@@ I Introduction
Quantum Error Correction (QEC) relies on the measurement of error syndromes to detect spurious evolutions of a physical system embedding quantum information [1]. For stabilizer codes [2], the regular approach consists in measuring a set of stabilizers and decoding the measurement record to apply corrective feedback or feedforward controls. Alternatively, QEC can be made autonomous by engineering adequate dissipation channels [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Autonomous QEC is only practical in the most simple decoding cases, in particular when each feedback control depends on the instantaneous value of a single stabilizer only. This is the case for most single-mode bosonic codes [18], in which a qubit encoded in an oscillator is stabilized by a few operators only. This approach was pioneered in the setting of two-legged cat qubits [8, 19, 20, 21, 22, 23, 24, 25]. The two-legged cat code manifold is spanned by two coherent states |±$|\pm\alpha\rangle$ of a target oscillator $a$. It is stabilized by a Lindblad dynamics entailed by a single dissipation channel $\dot{\bm{\rho}}_{a}=\gamma_{2}\mathcal{D}[{\bf a}^{2}-\alpha^{2}]\bm{\rho}_{a}$, where $\gamma_{2}$ is the dissipation rate and $\mathcal{D}[{\bf L}]\bm{\rho}={\bf L}\bm{\rho}{\bf L}^{\dagger}-\big{(}{\bf L}^{\dagger}{\bf L}\bm{\rho}+\bm{\rho}{\bf L}^{\dagger}{\bf L}\big{)}/2$. The Lindblad operator ${\bf a}^{2}-\alpha^{2}$ is of order two in the annihilation and creation operators $({\bf a},{\bf a^{\dagger}})$ and by extension, we characterize the stabilizing dissipation channel as order-two. It may be engineered by activating a two-to-one photon exchange interaction between the target oscillator and an auxiliary mode $b$, combined with a resonant drive on $b$ to yield the Hamiltonian ${\bf H}_{2}=g_{2}({\bf a}^{2}-\alpha^{2}){\bf b^{\dagger}}+\mathrm{h.c.}$ If the auxiliary mode dissipates excitations into a cold environment at rate $\kappa_{b}$ much larger than the two-to-one photon exchange rate $g_{2}$, the multimode dynamics reduces to the desired single-mode Lindblad dynamics of the target mode with $\gamma_{2}=4g_{2}^{2}/\kappa_{b}$ [26, 27] [^1] This hierarchy is needed to adiabatically eliminate the mode $b$ from the dynamics, but does not need to be respected for the cat manifold to be stabilized [27]. Unless the two mode resonance frequencies are precisely targeted to reach $\omega_{b}=2\omega_{a}$ [24], the interaction needs to be parametrically modulated at $2\omega_{a}-\omega_{b}$ to become resonant. This can be realized with a variety of Josephson circuits such as the one pictured in Fig. 1a, all relying on the four-wave mixing capabilities of Josephson junctions. This strategy was successfully deployed to robustly suppress bit-flips of two-legged cat qubits. On the down side, this code does not offer any protection against phase-flips entailed by single-photon loss in the target mode, which is the dominant noise channel in most experimental platforms.
@@
In order to generalize the approach to codes whose states have more complex structures in phase-space and offer some measure of protection against phase-flips, one typically needs higher order dissipation channels and thus higher order interactions. In particular, the order-four channel corresponding to the Lindblad dissipator $\gamma_{4}\mathcal{D}[{\bf a}^{4}-\alpha^{4}]$ stabilizes a manifold of four coherent states, from which one can single out a sub-manifold of given parity to define a four-legged cat qubit [8]. This qubit is protected against bit-flips as its two-legged counterpart, and single-photon loss events can be detected without perturbing the encoded information by monitoring the photon number parity [8, 29, 30, 31, 32], thus suppressing phase-flips at first-order. Note that even higher order dissipation channels can correct parity flips autonomously [33] or stabilize other rotation-symmetric codes [34] and translation-symmetric codes [35, 14, 16], offering a more robust protection of the encoded information. In earlier experimental works, high-order dissipation channels were obtained from low-order interactions. Indeed, autonomous parity recovery was achieved by combining dispersive interactions and a set of single and two-photon drives [33], autonomous stabilization of the GKP code was achieved by combining Rabi interactions and sequences of single-photon drives [14, 16, 36], and a four-to-two photon exchange interaction—which could be used to engineer a four-photon dissipation channel—was obtained by cascading off-resonant two-to-one photon exchange interactions in a Raman configuration [37] [^2] In that work, the auxiliary mode was not designed to dissipate excitations at a sufficiently large rate to turn the interaction into an effective four-photon dissipation channel.. Unfortunately, this approach can lead to the opening of spurious dissipation channels entailed by each low-order primitive involved in the generation of effective high-order interactions. In principle, the problem can be addressed using a biased-noise auxiliary mode [39, 40, 41], error-transparent interactions [42, 43, 44, 45, 25] or carefully controlling the noise spectrum of the dissipative environment [46], but the experimental implementation of these ideas is a major challenge.
@@
Figure 1: Dissipation engineering in a Josephson circuit. a). A target oscillator (coined the memory, blue) is coupled to an auxiliary dissipative mode (coined the buffer, orange) by a SQUID participating in both modes. When the circuit loops are biased with magnetic fluxes $\Phi_{J}$ and $\Phi_{L}$ following Eq. (1) with $\epsilon(t)$ a pump oscillating at $4\omega_{a}-\omega_{b}$, the SQUID mediates a Kerr-free four-to-one photon exchange interaction, opening a four-photon dissipation channel acting on the memory. The mode impedances $Z_{a,b}$ set the strength of the interaction and are of the order of $R_{q}/10=650~{}\Omega$ in our experiment. b) Energy level-diagram featuring the four-to-one photon pump (purple wriggled arrow) and the buffer single-photon dissipation (orange wriggled arrow). In presence of strong non-linearity, energy levels are unevenly spaced and the pump drives a single transition only (dashed purple arrow represents an off-resonantly driven transition). c) Schematic representation of the circuit used in our experiment. A three-pad superconducting structure hosts the memory and buffer modes with respective anti-symmetric and symmetric voltages across the pads. A distributed $\lambda/2$ resonator capacitively coupled to the structure forms a third mode employed for readout. Foster synthesis of the linear circuit connected to the SQUID yields an effective circuit as in (a) but for the presence of a third mode and for spurious inductances in the arms of the SQUID, responsible for residual non-linearities (see Fig. 2a). Magnetic biases are applied through two flux lines (top). A charge line (bottom) is capacitively connected to the circuit and enables linear drives and single-photon dissipation of the buffer. All lines are filtered by periodic loading to limit single-photon dissipation in the memory [47]. d) Optical microscope image of the sample, featuring all the structures pictured in (c) with matching false colors, surrounded by a ground plane (brown).
@@
An alternative strategy circumventing these pitfalls consists in directly engineering high-order dissipation channels from high-order interactions. In this work, we take this approach and aim at engineering a four-photon loss channel—a key primitive to stabilize four-legged cat qubits—through genuine high-order interactions. Adapting the methods developed to stabilize two-legged cat qubits [21], we proceed by activating a resonant four-to-one photon exchange interaction between a long-lived superconducting resonator (coined the memory) and an auxiliary dissipative mode (coined the buffer). In the idealized setting depicted in Fig. 1a, two $LC$ oscillators are connected in series with a Superconducting Quantum Interference Device (SQUID). The SQUID enables mixing of four waves from the memory $a$, one wave from the buffer $b$, and one pump wave providing the energy needed for the interaction to be resonant. In detail, the circuit features two superconducting loops which are biased with magnetic fields
$\Phi_{J}(t)=\Phi_{0}\big{(}\frac{1}{2}-\frac{\epsilon(t)}{\pi}\big{)}\qquad\qquad\Phi_{L}(t)=\Phi_{0}\frac{\epsilon(t)}{2\pi}$ (1)
where $\Phi_{0}=h/(2e)$ is the flux quantum. The AC signal $\epsilon$, referred to as the pump, is much smaller than 1. In this regime, we express the Josephson potential of the SQUID as a function of the reduced phase $\varphi$ across the modes as $V_{\mathrm{SQUID}}(\varphi)=-E_{J}\cos(\varphi+2\pi\frac{\Phi_{L}}{\Phi_{0}})-E_{J}\cos(\varphi+2\pi\frac{\Phi_{L}+\Phi_{J}}{\Phi_{0}})\simeq 2E_{J}\epsilon(t)\sin(\varphi)$. We note that the potential is $\Phi_{0}$-periodic in $\Phi_{J}$ and $\Phi_{L}$, and offsetting $\Phi_{L}$ by $\Phi_{0}/2$ flips its sign. Expressed in terms of the modes creation and annihilation operators, the phase operator reads $\bm{\varphi}=\varphi^{ZPF}_{a}({\bf a}+{\bf a}^{\dagger})+\varphi^{ZPF}_{b}({\bf b}+{\bf b}^{\dagger})$ where the vacuum fluctuations $\varphi^{ZPF}_{i}$ of each mode are set by the mode impedance following $\varphi^{ZPF}_{i}=(\pi\frac{Z_{i}}{R_{q}})^{\frac{1}{2}}$ and $R_{q}\simeq 6.5~{}\mathrm{k}{\Omega}$ is the resistance quantum [48, 49]. We can then write the circuit Hamiltonian
${\bf H}(t)=\hbar\omega_{a}{\bf a}^{\dagger}{\bf a}+\hbar\omega_{b}{\bf b}^{\dagger}{\bf b}+2E_{J}\epsilon(t)\sin(\bm{\varphi})$ (2)
When $\epsilon=0$, the Hamiltonian (2) represents two uncoupled harmonic oscillators. By letting $\epsilon(t)=\xi e^{i(4\omega_{a}-\omega_{b})t}+\mathrm{c.c}.$, the SQUID enables a resonant four-to-one photon exchange interaction between the memory and the buffer—stemming from the fifth order term in the sine expansion. In the rotating frame of each mode and applying the Rotating Wave Approximation (RWA), we get [47]
In this expression, :_::\_:: _ : denotes the normal ordering of an operator, $g_{4}=2\frac{E_{J}}{\hbar 4!}\xi e^{-(\varphi^{{ZPF}^{2}}_{a}+\varphi^{{ZPF}^{2}}_{b})/2}\varphi^{{ZPF}^{4}}_{a}\varphi^{ZPF}_{b}$ depends strongly on the phase fluctuations of each mode across the SQUID, and $P_{a}$ and $P_{b}$ are polynomials boiling down to unit factors if the SQUID sine potential is truncated to leading order [47]. Crucially, no other non-linear term in the expansion of ${\bf H}(t)$ survives the RWA when the mode frequencies are not commensurable. Therefore, the memory and buffer remain perfectly linear. In presence of strong single-photon dissipation at rate $\kappa_{b}\gg g_{4}$ in the buffer, the system dynamics reduces to that entailed by a four-photon dissipation channel $\frac{4g_{4}^{2}}{\kappa_{b}}\mathcal{D}[{\bf a}^{4}]$ on the memory [26].
When compared to the analogous case where the same circuit is employed to mediate a two-to-one photon exchange interaction [21, 22, 23, 25], several complications arise. First, the process we target involves the mixing of five waves from the memory and buffer, so that the vacuum fluctuations $\varphi^{ZPF}_{a}$, $\varphi^{ZPF}_{b}$ should not be much smaller than 1 for the interaction rate not to be negligible  [^3] We note that approximating the polynomials in Eq. (3) with unit factors is only valid when $|\varphi^{ZPF}_{a}({\bf a}+{\bf a^{\dagger}})|,|\varphi^{ZPF}_{b}({\bf b}+{\bf b^{\dagger}})|\ll 1$ so that, when the four-to-one interaction is employed to autonomously stabilize four-legged cat states with mean photon number $|\alpha|^{2}$, there exists a maximum value of $\varphi^{ZPF}_{a}\alpha$ beyond which the approximation is no longer valid. This can lead to distortion and eventually destabilization of the cat manifold.. Second, in this regime, any deviation from the idealized circuit considered so far may lead to significant spurious non-linearity of the memory. Indeed, irrespective of the exact mechanism breaking the circuit ideality, the leading non-linear term on the memory scales as $\varphi^{{ZPF}^{4}}_{a}$ (Kerr non-linearity) [47]. A consequence of strong non-linearity is that the memory energy levels are no longer evenly spaced, and the frequency of the four-to-one photon transitions $|n+4\rangle_{a}|0\rangle_{b}\leftrightarrow|n\rangle_{a}|1\rangle_{b}$ that one needs to address depends on the memory level index $n$ (see Fig. 1b). The dominant source of non-linearity in our system, which we investigate in Sec. II.3, involves a Cooper-pair pairing mechanism stemming from stray inductance of the SQUID tracks (purple inductor in Fig. 1c). Unfortunately, it is strong enough to prevent the pump from driving more that one specific four-to-one photon transition. Nevertheless, driving the transition $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$, we demonstrate that quartets of excitations can be dissipated at a rate an order of magnitude higher than the single-photon loss rate in our device, as detailed in the remaining of the paper. Given that stray inductances could be significantly reduced by simple design changes in a future iteration of the experiment, our results paves the way for the stabilization of a four-legged cat manifold by dissipation engineering.
@@ II.1 Design
The 2D superconducting circuit we design to implement the idealized circuit analyzed in the previous section is pictured in Fig. 1c-d. At its core lie three superconducting islands (blue and orange squares). The top and central islands are bridged by an Asymmetrically Threaded SQUID (ATS) [21] whose shunt superinductor is made of a chain of 25 Josephson junctions for a total nominal inductance of $L=28.5$ nH. A second identical inductor bridges the central and bottom islands. This symmetric structure supports the memory and buffer, with respective anti-symmetric and symmetric voltages across the pads. The SQUID participates strongly in both modes, resulting in strong phase-fluctuations as prescribed in the previous section. A distributed $\lambda/2$ resonator, colored in green and coined the readout mode, is capacitively coupled to the three-island structure with only a weak participation of the ATS. This mode was originally intended to provide a photon number resolving measurement of the memory leveraging the wave-mixing capabilities of the ATS [23]. In practice, the memory and readout modes were found to be dispersively coupled due to circuit non-idealities (see Sec. II.3) such that dispersive measurement was sufficient to characterize the state of the memory (see Sec. III). Performing Foster synthesis of the linear circuit connected to the SQUID [51, 48, 49], one derives the effective circuit pictured in Fig. 2a with the hierarchy $Z_{r}\ll Z_{a}\sim Z_{b}\sim R_{q}/10$. This effective circuit is analogous to the ideal circuit pictured in Fig. 1a but for the presence of a third mode and for the series inductances in the arms of the SQUID. Correspondence between fluxes applied through the ATS loops $(\Phi_{1},\Phi_{2})$ and fluxes biasing the ideal circuit is made through $\Phi_{L}=-\Phi_{1}$ and $\Phi_{J}=\Phi_{1}+\Phi_{2}$. The fluxes are induced by currents threaded through two dedicated feedlines, denoted as flux lines. We here note that AC fluxes are not independently calibrated in our experiment, such that the condition (1) on the ratio of AC fluxes is not necessarily satisfied. As a result, the modes frequency and non-linearity are slightly renormalized by the pump [47]. Finally, a third feedline, denoted as the charge line, is capacitively coupled to the three-pad structure and to the readout resonator. It allows direct linear drive of the modes, and sets the single-photon decay rate of the buffer and readout modes (see Table 1). All feedlines are filtered through sinusoidal modulation of their geometry to limit radiative decay of the memory [47].

### v2

@@ I Introduction
Quantum Error Correction (QEC) relies on the measurement of error syndromes to detect spurious evolutions of a physical system embedding quantum information [1]. For stabilizer codes [2], the regular approach consists in measuring a set of stabilizers and decoding the measurement record to apply corrective feedback or feedforward controls. Alternatively, QEC can be made autonomous by engineering adequate dissipation channels [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Autonomous QEC is only practical in the most simple decoding cases, in particular when each feedback control depends on the instantaneous value of a single stabilizer only. This is the case for most single-mode bosonic codes [22], in which a qubit encoded in an oscillator is stabilized by a few operators only. This approach was pioneered in the setting of two-legged cat qubits [10, 23, 24, 25, 26, 27, 28, 29]. The two-legged cat code manifold is spanned by two coherent states |±$|\pm\alpha\rangle$ of a target oscillator $a$. It is stabilized by a Lindblad dynamics entailed by a single dissipation channel $\dot{\bm{\rho}}_{a}=\gamma_{2}\mathcal{D}[{\bf a}^{2}-\alpha^{2}]\bm{\rho}_{a}$, where $\gamma_{2}$ is the dissipation rate and $\mathcal{D}[{\bf L}]\bm{\rho}={\bf L}\bm{\rho}{\bf L}^{\dagger}-\big{(}{\bf L}^{\dagger}{\bf L}\bm{\rho}+\bm{\rho}{\bf L}^{\dagger}{\bf L}\big{)}/2$. The Lindblad operator ${\bf a}^{2}-\alpha^{2}$ is of order two in the annihilation and creation operators $({\bf a},{\bf a^{\dagger}})$ and by extension, we characterize the stabilizing dissipation channel as second order. It may be engineered by activating a two-to-one photon exchange interaction between the target oscillator and an auxiliary mode $b$, combined with a resonant drive on $b$ to yield the Hamiltonian ${\bf H}_{2}=g_{2}({\bf a}^{2}-\alpha^{2}){\bf b^{\dagger}}+\mathrm{h.c.}$ If the auxiliary mode dissipates excitations into a cold environment at rate $\kappa_{b}$ much larger than the two-to-one photon exchange rate $g_{2}$, the multimode dynamics reduces to the desired single-mode Lindblad dynamics of the target mode with $\gamma_{2}=4g_{2}^{2}/\kappa_{b}$ [30, 31] [^1] This hierarchy is needed to adiabatically eliminate the mode $b$ from the dynamics, but does not need to be respected for the cat manifold to be stabilized [31]. Unless the two mode resonance frequencies are precisely targeted to reach $\omega_{b}=2\omega_{a}$ [28], the interaction needs to be parametrically modulated at $2\omega_{a}-\omega_{b}$ to become resonant. This can be realized with a variety of Josephson circuits such as the one pictured in Fig. 1a, all relying on the four-wave mixing capabilities of Josephson junctions. This strategy was successfully deployed to robustly suppress bit-flips of two-legged cat qubits. On the down side, this code does not offer any protection against phase-flips entailed by single-photon loss in the target mode, which is the dominant noise channel in most experimental platforms.
@@
In order to generalize the approach to codes whose states have more complex structures in phase-space and offer some measure of protection against phase-flips, one typically needs higher order dissipation channels and thus higher order interactions. In particular, the fourth order channel corresponding to the Lindblad dissipator $\gamma_{4}\mathcal{D}[{\bf a}^{4}-\alpha^{4}]$ stabilizes a manifold of four coherent states, from which one can single out a sub-manifold of given parity to define a four-legged cat qubit [10, 21]. This qubit is protected against bit-flips as its two-legged counterpart, and single-photon loss events can be detected without perturbing the encoded information by monitoring the photon number parity [10, 33, 34, 35, 36], thus suppressing phase-flips at first-order. Note that even higher order dissipation channels can correct parity flips autonomously [37] or stabilize other rotation-symmetric codes [38] and translation-symmetric codes [39, 17, 19], offering a more robust protection of the encoded information. In earlier experimental works, high-order dissipation channels were obtained from low-order interactions. As discussed in Sec. IV, this approach can lead to the opening of spurious dissipation channels directly corrupting the encoded information.
@@
Figure 1: Dissipation engineering in a Josephson circuit. a). A target oscillator (the memory, blue) is coupled to an auxiliary dissipative mode (the buffer, orange) by a SQUID participating in both modes. When the circuit loops are biased with magnetic fluxes $\Phi_{1}$ and $\Phi_{2}$ following Eq. (1) with $\epsilon(t)$ a pump oscillating at $4\omega_{a}-\omega_{b}$, the SQUID mediates a Kerr-free four-to-one photon exchange interaction, opening a four-photon dissipation channel acting on the memory. The mode impedances $Z_{a,b}$ set the strength of the interaction and are of the order of $R_{q}/10=650~{}\Omega$ in our experiment. b) Energy level-diagram featuring the four-to-one photon pump (purple wriggled arrow) and the buffer single-photon dissipation (orange wriggled arrow). In presence of strong non-linearity, energy levels are unevenly spaced and the pump drives a single transition only (dashed purple arrow represents an off-resonantly driven transition). c) Schematic representation of the circuit used in our experiment. A three-pad superconducting structure hosts the memory and buffer modes with respective anti-symmetric and symmetric voltages across the pads. A distributed $\lambda/2$ resonator capacitively coupled to the structure forms a third mode employed for readout. Foster synthesis of the linear circuit connected to the SQUID yields an effective circuit as in (a) but for the presence of a third mode and for series inductances in the arms of the SQUID, responsible for spurious non-linearities (see Fig. 2a and text). Magnetic biases are applied through two flux lines (top). A charge line (bottom) is capacitively connected to the circuit and enables linear drives and single-photon dissipation of the buffer. All lines are filtered by periodic loading to limit single-photon dissipation in the memory [40]. d) Optical microscope image of the sample, featuring all the structures pictured in (c) with matching false colors, surrounded by a ground plane (brown).
@@
An alternative strategy circumventing this pitfall consists in directly engineering high-order dissipation channels from high-order interactions. In this work, we take this approach and aim at engineering a four-photon loss channel—a key primitive to stabilize four-legged cat qubits—through genuine high-order interactions. Adapting the methods developed to stabilize two-legged cat qubits [25], we proceed by activating a resonant four-to-one photon exchange interaction between a long-lived superconducting resonator (coined the memory) and an auxiliary dissipative mode (coined the buffer). In the idealized setting depicted in Fig. 1a, two $LC$ oscillators are connected in series with a balanced Superconducting Quantum Interference Device (SQUID). The junctions of the SQUID enable mixing of four waves from the memory $a$, one wave from the buffer $b$, and one pump wave providing the energy needed for the interaction to be resonant.
In detail, the circuit features two superconducting loops which are biased with magnetic fields
$\Phi_{1}(t)=\varphi_{0}\big{(}\pi+\epsilon(t)\big{)}\qquad\qquad\Phi_{2}(t)=\varphi_{0}\epsilon(t)$ (1)
where $\varphi_{0}=\hbar/(2e)$ is the reduced flux quantum and the AC signal $\epsilon$ is referred to as the pump. The SQUID couples to the sum of the modes reduced phase operators $\bm{\varphi}=\varphi^{ZPF}_{a}({\bf a}+{\bf a}^{\dagger})+\varphi^{ZPF}_{b}({\bf b}+{\bf b}^{\dagger})$ via
${\bf H}_{J}(t)$ $=-E_{J}\cos(\bm{\varphi}-\frac{\Phi_{1}}{\varphi_{0}})-E_{J}\cos(\bm{\varphi}+\frac{\Phi_{2}}{\varphi_{0}})$ (2)
$=2E_{J}~{}\sin\big{(}\epsilon(t)\big{)}\sin(\bm{\varphi})$
Expanding the sine terms, we find that the SQUID enables the mixing of an odd number of pump waves with an odd number of mode waves. In particular, by letting $\epsilon(t)=\xi e^{i(4\omega_{a}-\omega_{b})t}+\mathrm{c.c}.$, we activate a resonant four-to-one photon exchange interaction between the memory and the buffer
${\bf H}_{4}=\hbar g_{4}{\bf a}^{4}{\bf b}^{\dagger}~{}+~{}\mathrm{h.c.}$ (3)
where the exchange rate $g_{4}\simeq 2\frac{E_{J}}{\hbar 4!}\xi\varphi^{{ZPF}^{4}}_{a}\varphi^{ZPF}_{b}$ depends strongly on the vacuum phase fluctuations of the modes. For each mode $i$, these are set by the mode impedance following $\varphi^{ZPF}_{i}=(\pi\frac{Z_{i}}{R_{q}})^{\frac{1}{2}}$ and $R_{q}\simeq 6.5~{}\mathrm{k}{\Omega}$ is the resistance quantum [41, 42, 43]. Therefore, we require the modes impedance to be large for the interaction rate to be significant (see Sec. IV for a discussion on the optimal parameter regime). In presence of strong single-photon dissipation at rate $\kappa_{b}\gg g_{4}$ in the buffer, the system dynamics reduces to that entailed by a four-photon dissipation channel $\frac{4g_{4}^{2}}{\kappa_{b}}\mathcal{D}[{\bf a}^{4}]$ on the memory [30].
Crucially, no other non-linear term in the expansion of ${\bf H}_{J}(t)$ survives the Rotating Wave Approximation (RWA) when the mode frequencies are not commensurable. Therefore, the memory and buffer remain in principle perfectly linear. However, in our experiment, stray inductance of the SQUID tracks (purple inductor in Fig. 1c) induce a spurious non-linear mechanism that is not canceled at the system working point (1). As a result, the memory is sufficiently anharmonic that the four-to-one photon transitions $|n+4\rangle_{a}|0\rangle_{b}\leftrightarrow|n\rangle_{a}|1\rangle_{b}$ (where $|n\rangle_{i}$ denotes a Fock state of mode $i$) cannot be indistinguishably driven with a single pump microwave (see Fig. 1b). Nevertheless, driving the transition $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$, we demonstrate that quartets of excitations can be dissipated at a rate an order of magnitude higher than the single-photon loss rate in our device, as detailed in the remaining of the paper. We review in Sec. IV simple design changes that should significantly suppress undesired non-linear mechanisms in a future iteration of the experiment, paving the way for the stabilization of the four-legged cat manifold.
@@ II.1 Design
The 2D superconducting circuit we design to implement the idealized circuit analyzed in the previous section is pictured in Fig. 1c-d. At its core lie three superconducting islands (blue and orange squares). The top and central islands are bridged by an Asymmetrically Threaded SQUID (ATS) [25] whose shunt superinductor is made of a chain of 25 Josephson junctions for a total nominal inductance of $L=28.5$ nH. A second identical inductor bridges the central and bottom islands. This symmetric structure supports the memory and buffer, with respective anti-symmetric and symmetric voltages across the pads. The SQUID participates strongly in both modes, resulting in strong phase-fluctuations as prescribed in the previous section. A distributed $\lambda/2$ resonator, colored in green and coined the readout mode, is capacitively coupled to the three-island structure with only a weak participation of the ATS. It is employed to probe the state of the memory as detailed in Sec.III.1.
To model and quantize our system, we perform Foster synthesis of the linear circuit connected to the SQUID [44, 41, 43], yielding the three system modes (labeled $a$, $b$, $r$) represented by a chain of parallel $LC$ oscillators with the hierarchy $Z_{r}\ll Z_{a}\sim Z_{b}\sim R_{q}/10$ (see Fig. 2a). The ATS shunt inductor is accounted for in this synthesis, but not the SQUID itself, whose arms each feature a junction and a small series inductance. The SQUID is subsequently connected in parallel with the chain to obtain the effective circuit pictured in Fig. 2a. It is analogous to the ideal circuit pictured in Fig. 1a but for the presence of a third mode and for the series inductances in the arms of the SQUID. The fluxes (1) biasing the loops of the ATS are delivered through two dedicated feedlines, denoted as flux lines 222 The fluxes $\Phi_{1}$ and $\Phi_{2}$ that thread the physical circuit loops can be formally allocated to each arm of the SQUID, such that they are the same fluxes threading the effective circuit loops.. A third feedline, denoted as the charge line, is capacitively coupled to the three-pad structure and to the readout resonator. It allows direct linear drive of the modes and sets the single-photon decay rate of the buffer and readout modes (see Table 1). Note that in principle, one could leverage the distinct symmetries of the memory and buffer modes to selectively couple a feedline to the buffer, thus preventing radiative decay of memory photons. This was not attempted in our experiment, in which the coupling of the feedline is of the same order for all modes. Instead, we filter all feedlines by sinusoidally modulating their geometry, thereby suppressing radiative decay at the memory frequency [40].